# GRU: Mitigating the Trade-off between Unlearning and Retention for Large Language Models

### v1

@@ 3 Gradient Rectified Unlearning
As discussed above, many methods have been developed to mitigate excessive unlearning. However, these achievements often result in an inevitable trade-off between removal and retention—improvements in maintaining the overall performance typically occur at the expense of weakened strength of unlearning. Considering that both goals of removal and reliable retention, are necessary to unlearn practical LLM, this trade-off is detrimental to practical LLM unlearning. If the model is compromised on removal, then privacy breaches and harmful behaviors may result; if the model is compromised on retention, it may adversely affect the model’s overall utility, adversely affecting its commercial value.
In this paper, rather than developing new methods that can better balance the trade-off between removal and retention, we turn our focus toward directly breaking this dichotomy. To put it another way, it is our intention to explore frameworks that are capable of delivering improved unlearning without compromising the overall utility of the model.
@@ 3.1 Motivation and The Proposed Framework
$\boldsymbol{\theta}^{(t+1)}\leftarrow\boldsymbol{\theta}^{(t)}-\texttt{lr}\cdot\boldsymbol{g}^{(t)}_{\mathrm{u}}.$ (6)
@@
Therein, lr denotes the (un) learning rate and $\boldsymbol{g}^{(t)}_{\mathrm{u}}=\nabla_{\boldsymbol{\theta}}\mathcal{L}(\tilde{\mathcal{D}}_{\mathrm{u}}^{(t)};\boldsymbol{\theta}^{(t)})$ with $\tilde{\mathcal{D}}^{(t)}_{\mathrm{u}}$ the mini-batch of size $b$ sampled from $\mathcal{D}_{\mathrm{u}}$ and $\mathcal{L}$ being any unlearning loss mentioned in Section 2, e.g., GA, GD, or NPO. This direct updating rule has proven to be unreliable in terms of retention, leading to the undesirable trade-off between retention and removal, which is widely mentioned in many previous works [11, 14, 16].
@@
This notable drawback motivates us to replace $\boldsymbol{g}^{(t)}_{\mathrm{u}}$ in Eq. (6) with its constrained version $\tilde{\boldsymbol{g}}_{\mathrm{u}}^{(t)}$: We incorporate the retain loss $\mathcal{R}$ as in Eq. (2), along with the corresponding gradients ${\boldsymbol{g}}_{\mathrm{r}}^{(t)}=\nabla_{\boldsymbol{\theta}}\mathcal{R}(\mathcal{D}_{\mathrm{r}};\boldsymbol{\theta}^{(t)})$, typically estimated by random mini-batch drawn from $\mathcal{D}_{\mathrm{r}}$. Then, we assert that the adjusted gradients $\tilde{\boldsymbol{g}}_{\mathrm{u}}^{(t)}$ should meet the condition as
$\operatorname*{arg\,min}_{\tilde{\boldsymbol{g}}_{\mathrm{u}}^{(t)}}\hskip 2.0pt$ $\|\tilde{\boldsymbol{g}}_{\mathrm{u}}^{(t)}-\boldsymbol{g}_{\mathrm{u}}^{(t)}\|^{2}$ (7)
s.t. $\langle\tilde{\boldsymbol{g}}_{\mathrm{u}}^{(t)},{\boldsymbol{g}}_{\mathrm{r}}^{(t)}\rangle\geq 0.$
@@
The objective $\min\|\tilde{\boldsymbol{g}}_{\mathrm{u}}^{(t)}-\boldsymbol{g}_{\mathrm{u}}^{(t)}\|^{2}$ ensures that the constrained gradients remain close to their original values. Meanwhile, the constraint $\langle\tilde{\boldsymbol{g}}_{\mathrm{u}}^{(t)},{\boldsymbol{g}}_{\mathrm{r}}^{(t)}\rangle\geq 0$ guarantees that there will be no adverse effect on model performance in retain data after updates. Overall, Eq. (7) encapsulates our principle that the removal of targeted knowledge should occur under strict conditions that ensure the retention performance on non-targeted data, thereby mitigating the inherent trade-off associated with such an approach. It is our intention to mitigate the trade-off by adjusting the gradient direction, hence we name the proposed unlearning framework as gradient rectified unlearning (GRU).
@@
The rationale behind $\langle\tilde{\boldsymbol{g}}_{\mathrm{u}}^{(t)},{\boldsymbol{g}}_{\mathrm{r}}^{(t)}\rangle\geq 0$ for retention is simple: Assuming the model is locally linear [26], we can approximate the expected loss change for the retain data as $\mathcal{R}(\boldsymbol{\theta}+\texttt{lr}\cdot\tilde{\boldsymbol{g}}_{\mathrm{u}})-\mathcal{R}(\boldsymbol{\theta})\approx\texttt{lr}\langle\tilde{\boldsymbol{g}}_{\mathrm{u}}^{(t)},{\boldsymbol{g}}_{\mathrm{r}}^{(t)}\rangle$. As observed, a positive $\langle\tilde{\boldsymbol{g}}_{\mathrm{u}}^{(t)},{\boldsymbol{g}}_{\mathrm{r}}^{(t)}\rangle$ implies that the loss $\mathcal{R}$ does not deteriorate following gradient updates, thereby ensuring the goal of retention. Later, we will show that the condition expressed in Eq. (7) remains valid under some less stringent assumptions, demonstrating its practical applicability and highlighting the validity of the method.
@@ 3.2 Realizations
Figure 5: Illustration of GRU Updating Rule. Panels (a)-(b) display situations where the angles between the gradient vectors (red dashed and blue dashed arrows) are obtuse, violating the constraint in Eq. (7). In these cases, the gradients should be projected onto the orthogonal plane of the conflicting retain gradient. Panel (c) illustrates a scenario with an acute angle between the gradient vectors, adhering to the constraint in Eq. (7) and falling within the retention-safe feasible region (red half-circle), thus requiring no further adjustment.
Throughout this section we will explore details on how to implement GRU, focusing on the closed-form solution that GRU provides as well as additional strategies to enhance its reliability in the field.
@@
$\tilde{\boldsymbol{g}}_{\mathrm{u}}^{(t)}=$ $\boldsymbol{g}_{\mathrm{u}}^{(t)}+\frac{\max(-\langle\boldsymbol{g}_{\mathrm{u}}^{(t)},\boldsymbol{g}_{\mathrm{r}}^{(t)}\rangle,0)}{\|\boldsymbol{g}_{\mathrm{r}}^{(t)}\|^{2}}\boldsymbol{g}_{\mathrm{r}}^{(t)}$ (8)
$=$ $\boldsymbol{g}_{\mathrm{u}}^{(t)}+\frac{\|\boldsymbol{g}_{\mathrm{u}}^{(t)}\|\max(-\cos(\boldsymbol{g}_{\mathrm{u}}^{(t)},\boldsymbol{g}_{\mathrm{r}}^{(t)}),0)}{\|\boldsymbol{g}_{\mathrm{r}}^{(t)}\|}\boldsymbol{g}_{\mathrm{r}}^{(t)},$
where $\cos(\boldsymbol{g}_{\mathrm{u}}^{(t)},\boldsymbol{g}_{\mathrm{r}}^{(t)})={\langle\boldsymbol{g}_{\mathrm{u}}^{(t)},\boldsymbol{g}_{\mathrm{r}}^{(t)}\rangle}/({\|\boldsymbol{g}_{\mathrm{u}}^{(t)}\|\cdot\|\boldsymbol{g}_{\mathrm{r}}^{(t)}\|})$. For more detailed derivations, please refer to Appendix A.1.
Eq. (8) conveys a clear geometric interpretation, adjusting the original gradients $\boldsymbol{g}_{\mathrm{u}}^{(t)}$ onto the half-space defined by the constraint $\langle\boldsymbol{g}_{\mathrm{u}}^{(t)},\boldsymbol{g}_{\mathrm{r}}^{(t)}\rangle\geq 0$. If this constraint is already satisfied, then $\boldsymbol{g}_{\mathrm{u}}^{(t)}$ remains unchanged. Otherwise, $\boldsymbol{g}_{\mathrm{u}}^{(t)}$ will be projected in the direction that is orthogonal to $\boldsymbol{g}_{\mathrm{r}}^{(t)}$. Please refer to Figure 5 for some visual illustrations. Moreover, we present the implementation of our GRU in Algorithm 1, further elaborating on several key details as follows.
@@
1:Input: Initial parameters $\boldsymbol{\theta}_{\mathrm{org}}$, learning rate lr, number of iterations $T$, and hyperparameters $\gamma$, $\tau$.
2:Initialize the moving-average retention gradient: $\bar{\boldsymbol{g}}_{\mathrm{r}}^{(0)}\leftarrow\mathbf{0}$.
@@
4:     // 1) Sample a mini-batch from the unlearning dataset $\mathcal{D}_{u}$.
5:     $\mathcal{B}_{\mathrm{u}}^{(t)}\subset\mathcal{D}_{\mathrm{u}}$.
6:     // 2) Sample a mini-batch from the retention dataset $\mathcal{D}_{r}$.
7:     $\mathcal{B}_{\mathrm{r}}^{(t)}\subset\mathcal{D}_{\mathrm{r}}$.
8:     // 3) Compute the unlearning gradient $\boldsymbol{g}_{\mathrm{u}}$.
9:     $\boldsymbol{g}_{\mathrm{u}}^{(t)}\leftarrow\nabla_{\boldsymbol{\theta}}\,\mathcal{L}(\mathcal{B}_{\mathrm{u}}^{(t)};\boldsymbol{\theta}^{(t)})$.
10:     // 4) Compute the retention gradient $\boldsymbol{g}_{\mathrm{r}}$.
11:     $\boldsymbol{g}_{\mathrm{r}}^{(t)}\leftarrow\nabla_{\boldsymbol{\theta}}\,\mathcal{R}_{\mathrm{r}}(\mathcal{B}_{\mathrm{r}}^{(t)};\boldsymbol{\theta}^{(t)})$.
12:     // 5) Apply exponential moving average to stabilize the retention gradient.
13:     $\bar{\boldsymbol{g}}_{\mathrm{r}}^{(t)}\leftarrow(1-\gamma)\,\bar{\boldsymbol{g}}_{\mathrm{r}}^{(t-1)}\;+\;\gamma\,\boldsymbol{g}_{\mathrm{r}}^{(t)}$.
14:     // 6) Check whether gradients conflict: if the dot product is negative, project out the conflicting component.
15:     if  $\langle\boldsymbol{g}_{\mathrm{u}}^{(t)},\,\bar{\boldsymbol{g}}_{\mathrm{r}}^{(t)}\rangle<0$  then
16:         $\tilde{\boldsymbol{g}}_{\mathrm{u}}^{(t)}\leftarrow\boldsymbol{g}_{\mathrm{u}}^{(t)}\;-\;\frac{\langle\boldsymbol{g}_{\mathrm{u}}^{(t)},\,\bar{\boldsymbol{g}}_{\mathrm{r}}^{(t)}\rangle}{\|\bar{\boldsymbol{g}}_{\mathrm{r}}^{(t)}\|^{2}}\,\bar{\boldsymbol{g}}_{\mathrm{r}}^{(t)}$.
17:     else
18:         $\tilde{\boldsymbol{g}}_{\mathrm{u}}^{(t)}\leftarrow\boldsymbol{g}_{\mathrm{u}}^{(t)}$.
19:     end if
20:     // 7) Optionally clip the rectified gradient norm to avoid overly large updates.
21:     if  $\|\tilde{\boldsymbol{g}}_{\mathrm{u}}^{(t)}\|\;>\;\tau$  then
22:         $\tilde{\boldsymbol{g}}_{\mathrm{u}}^{(t)}\leftarrow\tau\,\frac{\tilde{\boldsymbol{g}}_{\mathrm{u}}^{(t)}}{\|\tilde{\boldsymbol{g}}_{\mathrm{u}}^{(t)}\|}$.
23:     end if
24:     // 8) Update the parameters using the rectified gradient.
25:     $\boldsymbol{\theta}^{(t+1)}\leftarrow\boldsymbol{\theta}^{(t)}-\,\texttt{lr}^{(t)}\>\tilde{\boldsymbol{g}}_{\mathrm{u}}^{(t)}$;
26:end for
27:Return $\boldsymbol{\theta}^{(T)}$.
@@
Stable Estimation. In practice, we typically utilize stochastic mini-batches of data to estimate the exact values of ${\boldsymbol{g}}_{\mathrm{r}}^{(t)}$ outlined in Eq. (7), which may lead to stochastic errors especially when the batch size is small, often the cases for LLM unlearning. Therefore, we employ the exponential moving average (EMA) to mitigate the additional computation costs associated with increasing batch sizes, namely,
$\bar{\boldsymbol{g}}_{\mathrm{r}}^{(t)}=(1-\gamma)\bar{\boldsymbol{g}}_{\mathrm{r}}^{(t-1)}+\gamma{\boldsymbol{g}}_{\mathrm{r}}^{(t)},$ (9)
@@
where $\gamma\in[0,1)$ is the smoothing parameter, of which smaller values suggest that a broader range of recent batches is covered, indicating a large batch size implicitly. It is worth noting that EMA is an approximation of using large batch sizes, given that $\boldsymbol{\theta}$ itself is also updated throughout the steps $t$. Therefore, selecting the appropriate value for $\gamma$ is crucial, as it involves balancing the representation of a larger batch size against minimizing the induced errors.
@@
Gradient Clipping. Due to stochastic variations and low-order approximations, the rectified gradients may inadvertently encroach upon regions that may decrease retention. To further enhance the practical reliability of our GRU, we further constrain the gradient norm via gradient clipping, following many previous works such as [16, 26]. Specifically, the gradients are scaled down to ensure it stays within a bounded range, i.e.,
$\tilde{\boldsymbol{g}}_{\mathrm{u}}^{(t)}\leftarrow\begin{cases}\tilde{\boldsymbol{g}}_{\mathrm{u}}^{(t)},&\text{if }\|\tilde{\boldsymbol{g}}_{\mathrm{u}}^{(t)}\|\leq\tau\\ \tau{\tilde{\boldsymbol{g}}_{\mathrm{u}}^{(t)}}/{\|\tilde{\boldsymbol{g}}_{\mathrm{u}}^{(t)}\|},&\text{if }\|\tilde{\boldsymbol{g}}_{\mathrm{u}}^{(t)}\|>\tau\end{cases}\end{aligned},$ (10)

### v3

@@ 3 Gradient Rectified Unlearning
As discussed above, many methods have been developed to mitigate excessive unlearning. However, these achievements often result in an inevitable trade-off between removal and retention—improvements in maintaining the overall performance typically occur at the expense of weakened strength of unlearning. This trade-off is detrimental to practical LLM unlearning, since both the goals of removal and reliable retention are essential: Compromising on removal risks privacy breaches and harmful behaviors; compromising on retention can adversely affect the overall utility of the model, negatively affecting its commercial value.
In this paper, rather than developing new methods that can better balance the trade-off between removal and retention, we turn our focus toward directly breaking this dichotomy. In other words, we aim to explore frameworks in which improved unlearning does not compromise the overall utility.
@@ 3.1 Motivation and The Proposed Framework
$\bm{\theta}^{(t+1)}\leftarrow\bm{\theta}^{(t)}-\texttt{lr}\cdot\bm{g}^{(t)}_{\mathrm{u}}.$ (6)
@@
Therein, lr denotes the (un) learning rate and $\bm{g}^{(t)}_{\mathrm{u}}=\nabla_{\bm{\theta}}\mathcal{L}(\tilde{\mathcal{D}}_{\mathrm{u}}^{(t)};\bm{\theta}^{(t)})$ with $\tilde{\mathcal{D}}^{(t)}_{\mathrm{u}}$ the mini-batch of size $b$ sampled from $\mathcal{D}_{\mathrm{u}}$ and $\mathcal{L}$ being any unlearning loss mentioned in Section 2, e.g., GA, GD, or NPO. This direct updating rule has proven to be unreliable in terms of retention, leading to the undesirable trade-off between retention and removal, which is widely mentioned in many previous works (Wang et al., 2024a; Liu et al., 2024b; Maini et al., 2024).
@@
This notable drawback motivates us to replace $\bm{g}^{(t)}_{\mathrm{u}}$ in Eq. (6) with its constrained version $\tilde{\bm{g}}_{\mathrm{u}}^{(t)}$: We incorporate the retain loss $\mathcal{R}$ as in Eq. (2), along with the corresponding gradients ${\bm{g}}_{\mathrm{r}}^{(t)}=\nabla_{\bm{\theta}}\mathcal{R}(\mathcal{D}_{\mathrm{r}};\bm{\theta}^{(t)})$, typically estimated by random mini-batch drawn from $\mathcal{D}_{\mathrm{r}}$. Then, we assert that the adjusted gradients $\tilde{\bm{g}}_{\mathrm{u}}^{(t)}$ should meet the condition as
$\operatorname*{arg\,min}_{\tilde{\bm{g}}_{\mathrm{u}}^{(t)}}\hskip 2.0pt$ $\|\tilde{\bm{g}}_{\mathrm{u}}^{(t)}-\bm{g}_{\mathrm{u}}^{(t)}\|^{2}$ (7)
s.t. $\langle\tilde{\bm{g}}_{\mathrm{u}}^{(t)},{\bm{g}}_{\mathrm{r}}^{(t)}\rangle\geq 0.$
@@
The objective $\min\|\tilde{\bm{g}}_{\mathrm{u}}^{(t)}-\bm{g}_{\mathrm{u}}^{(t)}\|^{2}$ ensures that the constrained gradients remain close to their original values. Meanwhile, the constraint $\langle\tilde{\bm{g}}_{\mathrm{u}}^{(t)},{\bm{g}}_{\mathrm{r}}^{(t)}\rangle\geq 0$ guarantees that the updates will not impair the model performance on retain data. Overall, Eq. (7) encapsulates our principle that the removal of targeted knowledge should occur under strict conditions that ensure the retention of performance on non-targeted data, thereby mitigating the inherent trade-off. As we mitigate the trade-off by adjusting the gradient direction, we name the corresponding unlearning framework as gradient rectified unlearning (GRU).
@@
The rationale behind $\langle\tilde{\bm{g}}_{\mathrm{u}}^{(t)},{\bm{g}}_{\mathrm{r}}^{(t)}\rangle\geq 0$ for retention is simple: Assuming the model is locally linear (Wortsman et al., 2022), we can approximate the expected loss change for the retain data as $\mathcal{R}(\bm{\theta}+\texttt{lr}\cdot\tilde{\bm{g}}_{\mathrm{u}})-\mathcal{R}(\bm{\theta})\approx-\texttt{lr}\langle\tilde{\bm{g}}_{\mathrm{u}}^{(t)},{\bm{g}}_{\mathrm{r}}^{(t)}\rangle$. As observed, a positive $\langle\tilde{\bm{g}}_{\mathrm{u}}^{(t)},{\bm{g}}_{\mathrm{r}}^{(t)}\rangle$ implies that the loss $\mathcal{R}$ does not deteriorate following gradient updates, thereby ensuring the goal of retention. Later, we will show that the condition expressed in Eq. (7) remains valid under some less stringent assumptions, further highlighting its practical applicability.
@@ 3.2 Realizations
Figure 5: Illustration of GRU Updating Rule. Panels (a)-(b) display situations where the angles between the gradient vectors (red dashed and blue dashed arrows) are obtuse, violating the constraint in Eq. (7). In these cases, the gradients should be adjusted orthogonally. Panel (c) illustrates a scenario with an acute angle between the gradient vectors, adhering to the constraint in Eq. (7) and falling within the retention-safe feasible region (red half-circle), thus requiring no further adjustment.
This section explores details to implement GRU, focusing on its closed-form solution as well as additional strategies to enhance its reliability in practice.
@@
$\tilde{\bm{g}}_{\mathrm{u}}^{(t)}=$ $\bm{g}_{\mathrm{u}}^{(t)}+\frac{\max(-\langle\bm{g}_{\mathrm{u}}^{(t)},\bm{g}_{\mathrm{r}}^{(t)}\rangle,0)}{\|\bm{g}_{\mathrm{r}}^{(t)}\|^{2}}\bm{g}_{\mathrm{r}}^{(t)}$ (8)
$=$ $\bm{g}_{\mathrm{u}}^{(t)}+\frac{\|\bm{g}_{\mathrm{u}}^{(t)}\|\max(-\cos(\bm{g}_{\mathrm{u}}^{(t)},\bm{g}_{\mathrm{r}}^{(t)}),0)}{\|\bm{g}_{\mathrm{r}}^{(t)}\|}\bm{g}_{\mathrm{r}}^{(t)},$
where $\cos(\bm{g}_{\mathrm{u}}^{(t)},\bm{g}_{\mathrm{r}}^{(t)})={\langle\bm{g}_{\mathrm{u}}^{(t)},\bm{g}_{\mathrm{r}}^{(t)}\rangle}/({\|\bm{g}_{\mathrm{u}}^{(t)}\|\cdot\|\bm{g}_{\mathrm{r}}^{(t)}\|})$. For more detailed derivations, please refer to Appendix A.1.
Eq. (8) conveys a clear geometric interpretation, adjusting the original gradients $\bm{g}_{\mathrm{u}}^{(t)}$ onto the half-space defined by the constraint $\langle\bm{g}_{\mathrm{u}}^{(t)},\bm{g}_{\mathrm{r}}^{(t)}\rangle\geq 0$. If this constraint is already satisfied, then $\bm{g}_{\mathrm{u}}^{(t)}$ remains unchanged. Otherwise, $\bm{g}_{\mathrm{u}}^{(t)}$ will be projected in the direction that is orthogonal to $\bm{g}_{\mathrm{r}}^{(t)}$. Please refer to Figure 5 for some visual illustrations. Moreover, we present the implementation of our GRU in Algorithm 1, further elaborating on several key details as follows.
@@
1:  Input: Initial parameters $\bm{\theta}_{\mathrm{org}}$, learning rate lr, number of iterations $T$, and hyperparameters $\gamma$, $\tau$.
2:  Initialize $\bar{\bm{g}}_{\mathrm{r}}^{(0)}=\mathbf{0}$;
@@
4:     sample the mini-batches of ${\mathcal{B}}_{\mathrm{u}}^{(t)}$ and ${\mathcal{B}}_{\mathrm{r}}^{(t)}$ from ${\mathcal{D}}_{\mathrm{u}}$ and ${\mathcal{D}}_{\mathrm{r}}$, respectively;
5:     ${\bm{g}}^{(t)}_{\mathrm{u}}\leftarrow\nabla_{\bm{\theta}}\mathcal{L}({\mathcal{B}}_{\mathrm{u}}^{(t)};\bm{\theta}^{(t)})$;
6:     ${\bm{g}}_{\mathrm{r}}^{(t)}\leftarrow\nabla_{\bm{\theta}}\mathcal{R}_{\mathrm{r}}({\mathcal{B}}_{\mathrm{r}}^{(t)};\bm{\theta}^{(t)});$
7:     $\bar{\bm{g}}_{\mathrm{r}}^{(t)}\leftarrow(1-\gamma)\bar{\bm{g}}_{\mathrm{r}}^{(t-1)}+\gamma{\bm{g}}_{\mathrm{r}}^{(t)}$;
8:     if  $\langle{\bm{g}}_{\mathrm{u}}^{(t)},\,\bar{\bm{g}}_{\mathrm{r}}^{(t)}\rangle<0$  then
9:        $\tilde{\bm{g}}_{\mathrm{u}}^{(t)}\leftarrow{\bm{g}}_{\mathrm{u}}^{(t)}-\frac{\langle{\bm{g}}_{\mathrm{u}}^{(t)},\,\bar{\bm{g}}_{\mathrm{r}}^{(t)}\rangle}{\|\bar{\bm{g}}_{\mathrm{r}}^{(t)})\|^{2}}\,\bar{\bm{g}}_{\mathrm{r}}^{(t)}$;
10:     else
11:        $\tilde{\bm{g}}_{\mathrm{u}}^{(t)}\leftarrow{\bm{g}}_{\mathrm{u}}^{(t)}$;
12:     end if
13:     if  $\|\tilde{\bm{g}}_{\mathrm{u}}^{(t)}\|>\tau$  then
14:        $\tilde{\bm{g}}_{\mathrm{u}}^{(t)}\leftarrow\tau{\tilde{\bm{g}}_{\mathrm{u}}^{(t)}}/{\|\tilde{\bm{g}}_{\mathrm{u}}^{(t)}\|}$;
15:     end if
16:     $\bm{\theta}^{(t+1)}\leftarrow\bm{\theta}^{(t)}-\texttt{lr}^{(t)}\tilde{\bm{g}}_{\mathrm{u}}^{(t)}$;
17:  end for
18:   Return $\bm{\theta}^{(T)}$.
@@
Stable Estimation. In practice, we typically utilize stochastic mini-batches of data to estimate the exact values of ${\bm{g}}_{\mathrm{r}}^{(t)}$ outlined in Eq. (7). Specifically, as shown in Algorithm 1, the mini-batches ${\mathcal{B}}_{\mathrm{u}}^{(t)}$ and ${\mathcal{B}}_{\mathrm{r}}^{(t)}$ serve as substitutes for the complete datasets $\mathcal{D}_{\mathrm{u}}$ and $\mathcal{D}_{\mathrm{r}}$. However, this may introduce stochastic errors, particularly when the batch size is small, which is commonly the case in LLM unlearning. Therefore, we employ the exponential moving average (EMA) to mitigate the additional computation costs associated with increasing batch sizes, namely,
$\bar{\bm{g}}_{\mathrm{r}}^{(t)}=(1-\gamma)\bar{\bm{g}}_{\mathrm{r}}^{(t-1)}+\gamma{\bm{g}}_{\mathrm{r}}^{(t)},$ (9)
@@
where $\gamma\in[0,1)$ is the smoothing parameter, with smaller values suggesting that a broader range of recent batches is covered, indicating a large batch size implicitly. It is worth noting that EMA is an approximation of using large batch sizes, given that $\bm{\theta}$ itself is also updated throughout the steps $t$. Therefore, selecting the appropriate value for $\gamma$ is crucial, as it involves balancing the representation of a larger batch size against minimizing the induced errors.
@@
Gradient Clipping. Due to stochastic variations and low-order approximations, the rectified gradients may inadvertently encroach upon regions that may decrease retention. To further enhance the practical reliability of our GRU, we further constrain the gradient norm via gradient clipping, following many previous works such as (Wortsman et al., 2022; Wang et al., 2024a). Specifically, the gradients are scaled down to ensure it stays within a bounded range, i.e.,
$\tilde{\bm{g}}_{\mathrm{u}}^{(t)}\leftarrow\begin{cases}\tilde{\bm{g}}_{\mathrm{u}}^{(t)},&\text{if }\|\tilde{\bm{g}}_{\mathrm{u}}^{(t)}\|\leq\tau\\ \tau{\tilde{\bm{g}}_{\mathrm{u}}^{(t)}}/{\|\tilde{\bm{g}}_{\mathrm{u}}^{(t)}\|},&\text{if }\|\tilde{\bm{g}}_{\mathrm{u}}^{(t)}\|>\tau\end{cases}\end{aligned},$ (10)